# Exploring Holomorphic Retracts

### v1

@@ 2.5. Fixed point sets and Retracts
Note that $F_{i}(z)$ is a holomorphic function from $\Delta^{N}$ to $\Delta$, because by condition (ii) of the theorem. Next we define $F(z):=(F_{1}(z),\ldots,F_{N}(z))$. $F$ is a holomorphic retraction map on $\Delta^{N}$ whose image is $span\{v_{k}\}_{k=1}^{r}\cap\Delta^{N}$, where $v_{k}=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}$, which completes the proof.
We already saw in the introduction section, there exists a linear retract of $\Delta^{N}$ which is not spanned by mutually disjoint supported vectors. The converse implication however is true and can now be derived as a corollary.
@@ Conclusion
The foregoing remarks for a product of Euclidean balls apply as well to the case of products of balls with polydiscs such as $\Delta^{k}\times\mathbb{B}$ (or a larger collection of Euclidean balls and polydiscs in any order) which are also homogeneous domains with all of is automorphisms being fractional-linear maps. We conclude by noting that in these cases the characterization of retracts provided by theorems 1.12 and 3.16, is complete.
@@
Let us begin by noting that it follows from theorem 1.10 proved in the foregoing section that every retract of a bounded balanced homogeneous domain is biholomorphic to the graph of some holomorphic function over a linear subspace of $D$.
Since $\Omega$ is homogenous, $\Omega/Aut(\Omega)$ is singleton and hence compact. By the Main theorem in [Fost], there exists a closed complex submanifold $Z$ of $\Omega$ such that M is biholomorphic to a locally trivial holomorphic fibre bundle with fibre $\mathbb{C}$ over $Z$. As mentioned in the begining, by theorem 1.10 we get that Z is biholomorphic to a linear retract $\Omega_{L}:=\Omega\cap L$ for some linear subspace $L$ of $\mathbb{C}^{N}$. We can cover $M$ by coordinate charts biholomorphic to $\{U_{\alpha}\times\mathbb{C}\}$ where $U_{\alpha}$ is an open cover of $Z$. The coordinate transformations from the $\alpha$ to the $\beta$- coordinates are of the form
@@
for some holomorphic functions $k_{\alpha\beta},h_{\alpha\beta}$ on $U_{\alpha}\cap U_{\beta},h_{\alpha\beta}\neq 0$. The data $h_{\alpha\beta}$ determine a multiplicative Cousin problem on $Z$. Since $Z$ is homeomorphic to polydisk and Oka’s theorem in [Oka](exposited in theorem 3.9 of [oka_0] as well), multiplicative Cousin problem for $Z$ can be solved. Hence we might have assumed already above that $h_{\alpha\beta}\equiv 1$ for all $\alpha,\beta$. However in that case, the data $k_{\alpha\beta}$ determine an additive Cousin problem. By Oka’s theorem in [Oka_1] (exposited in theorem 3.6 of [oka_0]), additive Cousin problem can be solved. Hence $M$ is biholomorphic to $Z\times\mathbb{C}$.
@@
We recall for one last time, that the class of homogeneous bounded balanced domains is precisely the class of bounded balanced symmetric domains and that any domain of this class, is biholomorphic to a Cartesian product of Cartan domains of types I – IV (cf. the second chapter, specifically remark 2.3.8 of the book [Jrncki_invrnt_dst] and references therein). It follows that a domain $\Omega$ as in the above theorem is a polyball (and convex), in-particular.
@@
It must first be noted that by Theorem 1.3, every holomorphic retract through the origin is actually linear i.e., $Y_{q}$ is a holomorphic retract of $D_{q}$ through the origin, then there exists a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$. Of-course, we are only interested in non-trivial retracts, so $Y_{q}$ is of the form $Y_{q}=D_{q}\cap Y$ for some $r$-dimensional $\mathbb{C}$-linear subspace $Y$ of $\mathbb{C}^{N}$. We only need to show that if $Y$ cannot be of the form $\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:z_{i}=a_{1,i}z_{1}+\ldots+a_{r,i}z_{r},~{}r+1\leq i\leq N\}$ of $\mathbb{C}^{N}$, where not all $(a_{i,j})$ are zero.
@@
To prove this proposition by contradiction, we assume that $Y_{q}$ is a linear retract of $D_{q}$ and a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$. Let $\{e_{i}:1\leq i\leq N\}$ be the standard basis of $\mathbb{C}^{N}$. Define another basis $\{v_{i}:1\leq i\leq N\}$ of $\mathbb{C}^{N}$, as follows:
@@
$t_{j}=\frac{1}{\left(1+\sum_{k=r+1}^{N}|a_{j,k}|^{q}\right)^{1/q}}$
@@
$e_{j}=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}v_{l}\right)$
After applying $P$ on both sides, we get
$P(e_{j})=\frac{1}{t_{j}}P(v_{j})-\left(\sum_{l=r+1}^{N}a_{j,l}P(v_{l})\right)$
@@
(5.1) $P(v_{j})=\left(u_{1,j},\ldots,u_{r,j},\sum_{k=1}^{r}a_{k,r+1}u_{k,j},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,j}\right)~{}\text{for}~{}r+1\leq j\leq N$
@@
$P(e_{j})=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}\left(u_{1,l},\ldots,u_{r,l},\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,l}\right)\right)\\ =\left(-\sum_{l=r+1}^{N}a_{j,l}u_{1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{j-1,l},1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l},-\sum_{l=r+1}^{N}a_{j,l}u_{j+1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{r,l},\right.\\ \left.a_{j,r+1}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,a_{j,N}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,N}u_{k,l}\right)$
@@
(5.2) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}\left|a_{j,s}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,s}u_{k,l}\right|^{q}\leq 1$
The following inequality holds for $r+1\leq j\leq N$:
(5.3) $\sum_{k=1}^{r}|u_{k,j}|^{q}+\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}~{}u_{k,j}\right|^{q}\leq 1$
By reverse triangle inequality for square roots: $|z-w|^{q}\geq\left||z|^{q}-|w|^{q}\right|$ for all $z,w\in\mathbb{C}$, inequality (5.2) reduces to
$-\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0$
@@
$-\sum_{l=r+1}^{N}|a_{j,l}|^{q}\left(1-\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}-\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}u_{k,j}\right|^{q}\right)+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}+\\ \sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0$
Simplifing above inequality, we get
$\sum_{l=r+1}^{N}|a_{j,l}|^{q}\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}\leq 0$
From this inequality, we obtain
@@
From inequality (5.2), we get
(5.5) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|\leq t_{j},\quad 1\leq j\leq r$
Since not all $(a_{i,j})$ are zero, atleast one of $a_{i,j}$ is non-zero. If $a_{j_{0},l_{0}}\neq 0$ for some $j_{0},l_{0}$ then from (5.4) we obtain that $u_{j,l_{0}}=0$ for all $j$ with $j\neq j_{0}$. From (5.3), we get $|u_{j_{0},l_{0}}|\leq t_{j_{0}}$. We can conclude that for all $(j,l)$, we have $|u_{j,l}|\leq t_{j}$. Using (5.5), we get

### v6

@@ 2.5. Fixed point sets and Retracts
Note that $F_{i}(z)$ is a holomorphic function from $\Delta^{N}$ to $\Delta$, because by condition (ii) of the theorem. Consider the map $F$ defined in terms of these $F_{i}$’s by $F(z):=(F_{1}(z),\ldots,F_{N}(z))$. It can be readily checked that $F$ is a holomorphic retraction map on $\Delta^{N}$ whose image is $span\{v_{k}\}_{k=1}^{r}\cap\Delta^{N}$, where $v_{k}=e_{j_{k}}+\sum_{i\notin J}c_{i,j_{k}}e_{i}$, which completes the proof.
As noted above, there exists a linear retract of $\Delta^{N}$ which is not spanned by mutually disjoint supported vectors. The converse implication however is true and can now be derived as a corollary.
@@ Conclusion
The foregoing remarks for a product of Euclidean balls apply as well to the case of products of balls with polydiscs such as $\Delta^{k}\times\mathbb{B}$ (or a larger collection of Euclidean balls and polydiscs in any order) which are also homogeneous domains with all of is automorphisms being fractional-linear maps. We conclude by noting that in these cases the characterization of retracts provided by theorems 1.15 and 3.20, is complete.
@@
The proof follows ideas from the artice [Fost] of Fornaess–Sibony. Let us begin by noting that it follows from theorem 1.13 proved in the foregoing section that every retract of a bounded balanced homogeneous domain is biholomorphic to the graph of some holomorphic function over a linear subspace of $D$.
Since $\Omega$ is homogeneous, $\Omega/{\rm Aut}(\Omega)$ is singleton and hence compact. By the Main theorem in [Fost], there exists a closed complex submanifold $Z$ of $\Omega$ such that M is biholomorphic to a locally trivial holomorphic fibre bundle with fibre $\mathbb{C}$ over $Z$. As mentioned in the beginning, by theorem 1.13 we get that Z is biholomorphic to a linear retract $\Omega_{L}:=\Omega\cap L$ for some linear subspace $L$ of $\mathbb{C}^{N}$. We can cover $M$ by coordinate charts biholomorphic to $\{U_{\alpha}\times\mathbb{C}\}$ where $U_{\alpha}$ is an open cover of $Z$. The coordinate transformations from the $\alpha$ to the $\beta$- coordinates are of the form
@@
for some holomorphic functions $k_{\alpha\beta},h_{\alpha\beta}$ on $U_{\alpha}\cap U_{\beta},h_{\alpha\beta}\neq 0$. The data $h_{\alpha\beta}$ determine a multiplicative Cousin problem on $Z$. Since $Z$ is homeomorphic to a polydisc, it follows by Oka’s theorem (theorem 3.9 of [oka_0] as well) that the multiplicative Cousin problem for $Z$ can be solved. Hence we might have assumed already above that $h_{\alpha\beta}\equiv 1$ for all $\alpha,\beta$. However in that case, the data $k_{\alpha\beta}$ determine an additive Cousin problem. By Oka’s theorem in (theorem 3.6 of [oka_0]), the additive Cousin problem on $Z$ can also be solved. Hence $M$ is biholomorphic to $Z\times\mathbb{C}$.
@@
We recall for one last time, that the class of homogeneous bounded balanced domains is precisely the class of bounded balanced symmetric domains and that any domain of this class, is biholomorphic to a Cartesian product of Cartan domains of types I – IV (the second chapter, specifically remark 2.3.8 of the book [Jrncki_invrnt_dst] and references therein). It follows that a domain $\Omega$ as in the above theorem is a polyball (and convex), in-particular.
@@
Before discussing our result as in the title of this section on the retracts of $\ell_{q}$ balls for $10<q<$ 1 we first review the case $q\geq 1$. The main theorem about this case was already mentioned in theorem 1.7. However that formulation is highly coordinate dependent. Here we first re-formulate it in-terms of commutavity with standard projections; the proof is based on an idea from the work [Baronti-Pappini] of Baronti – Pappini.
Theorem 5.1.
$R$ is a $k$-dimensional retract of the $\ell^{p}$-ball in $\mathbb{C}^{N}$ if and only if one of the following holds:
  1. (1)
    There exists a norm-one projection which maps $\mathbb{B}_{p}$ onto $R$ which commutes with the standard orthogonal projection onto one of the $(N-k)$ or $(N-(k-1))$-dimensional subspace spanned by co-ordinate vectors.
  2. (2)
    There exist a finite sequence of projections $P_{k},P_{k-1},\ldots,P_{1}$ such that for each $r=1,2,\ldots,k$, $P_{r}$ is a projection of norm one which maps $\mathbb{C}^{N}$ onto an $r$-dimensional subspace and $P_{k}(\mathbb{B}_{p})=R$.
Proof.
It is easy to check that one implication of the theorem holds by using the definition of norm one projection. So we have only to prove the other implication. Assume that $R$ is a $k$-dimensional linear retract of $\mathbb{B}_{p}$. By theorem 2.a.4 in [Lindstraus], we obtain that $R=L\cap\mathbb{B}_{p}$ where $L$ is a $k$-dimensional linear subspace of $\mathbb{C}^{N}$ spanned by mutually disjointly supported vectors. Hence $L$ can be written as the solution set of $(N-k)$ linear equations, where each linear equation depends on atmost two variables. After applying Gaussian elimination and renaming variables, there is no loss of generality assume that above mentioned system of equations can be written as $AZ=0$, where $A$ is a $(N-k)\times N$ matrix $[B~{}B^{\prime}~{}B^{\prime\prime}]^{T}$ and $Z=[z_{1}~{}\ldots~{}z_{N}]^{T}$. The submatrices $B,B^{\prime},B^{\prime\prime}$ are of the following form:
  1. (i)
    $B=[B_{0}~{}B_{1}~{}\ldots B_{r-1}]^{T}$, where for each $i=0,1,\ldots,r-1$, $B_{i}=[b_{k_{i}+1}\ldots b_{k_{i+1}}]^{T}$ and $0=k_{0}<k_{1}<\ldots<k_{r}$. Futhermore for each $i$, $B_{i}$ satisfies the following properties:
    1. (a)
      For $j=k_{i}+1$ to $k_{i+1}-1$, $supp(b_{j})\cap supp(b_{j+1})\neq\phi$ and $|supp(b_{j})\cap supp(b_{j+1})|=1$.
    2. (b)
      For $i=0,1,\ldots,r-1$, if we define $supp(B_{i}):=\bigcup_{j=k_{i}+1}^{k_{i+1}}supp(b_{j})$ then $supp(B_{i})\cap supp(B_{j})=\phi$ for $i\neq j$.
  2. (ii)
    $B^{\prime}$ is of the form $[b_{1}^{\prime}~{}\ldots~{}b_{l}^{\prime}]^{T}$, where $supp(b_{i}^{\prime})\cap supp(b_{j}^{\prime})=\phi$ for $i,j=1,2,\ldots,l$ with $i\neq j$.
  3. (iii)
    $B^{\prime\prime}$ is of the form $[e_{s_{0}+1}~{}\ldots e_{s_{1}}]^{T}$. where $s_{0}=k_{r}+r+2l,s_{1}=N-k+r+l$ and $e_{i}$ denotes the standard basis of $\mathbb{C}^{N}$.
Note that $|supp(B)|=|\bigcup_{i=0}^{r-1}supp(B_{i})|=k_{r}+r$, $|supp(B^{\prime})|=2l$ and $|supp(B^{\prime\prime})|=s_{1}-s_{0}$. Since the number of equation is $(N-k)$, we have $k_{r}+l+(s_{1}-s_{0})=N-k$. Because of this equation, we take $s_{1}=N-k+r+l$. Hence, we obtain the total number of dependent variables in the system of equations $AZ=0$ is equal to
$|supp(B)|+|supp(B^{\prime})|+|supp(B^{\prime\prime})|$ $=$ $k_{r}+r+2l+(s_{1}-s_{0})$
$=$ $k_{r}+r+2l+(N-k)-(k_{r}+l)$
$=$ $N-k+r+l.$
First we want to construct mutually disjoint supported vectors $(v_{i})_{i=1}^{k}$ such that $(v_{i})$ forms a basis for $L$. For each $i=1,\ldots,r$, consider the system of equations $B_{i}Z_{i}=0$, where $Z_{i}=[z_{k_{i-1}+1}~{}\ldots~{}z_{k_{i}}]^{T}$. This solution space is spanned by a non-zero vector $w_{i}=(\mu_{k_{i-1}+1},\ldots\mu_{k_{i}})\in\mathbb{C}^{k_{i}-k_{i-1}}$. Since every two co-ordinates of $w_{i}$ satisfies a linear equation in two variables, every component of $w_{i}$ is non-zero. For each $i=1,\ldots,r$, define $v_{i}:=\mu_{k_{i-1}+1}e_{k_{i-1}+1}+\ldots+\mu_{k_{i}}e_{k_{i}}$, where $e_{i}$ denote the standard basis in $\mathbb{C}^{N}$. Note that $v_{i}\in L$ and $supp(v_{i})=k_{i}-k_{i-1}$. For $i=r+1$ to $r+l$, choose a non-zero vector $v_{i}$ which is orthogonal to $b^{\prime}_{i-r}$ and $supp(v_{i})=supp(b^{\prime}_{i-r})$. For $i=r+l+1$ to $k$, $v_{i}=e_{i+t}$, where $t=N-k$.
Define $u_{i}:=\frac{v_{i}}{\|v_{i}\|_{p}}$ for $i=1,2,\ldots,k$. Note that $(u_{i})_{i=1}^{k}$ forms a mutually disjointly supported vectors and a basis for $L$. Observe that $|supp(u_{i})|=2$ for $i=r+1$ to $r+l$ and $|supp(u_{i})|=1$ for $i=r+l+1$ to $k$. Using the idea in the proof of theorem 2.a.4 in [Lindstraus], we can construct projection $P$ of norm $1$ from given vectors $u_{1},\ldots,u_{k}$. Let $\sigma_{j}:=supp(u_{j})$ for $j=1,2,\ldots,k$. If we write $u_{j}=\sum_{i\in\sigma_{j}}\lambda_{i}e_{i}$, then we define the linear functional $u_{j}^{*}=\sum_{i\in\sigma_{j}}\frac{|\lambda_{i}|^{p}}{\lambda_{i}}e_{i}^{*}$. Observe that for $i=1,2,\ldots,r$, the linear functional $u_{i}^{*}$ depends only on the variables $z_{k_{i-1}+1},\ldots,z_{k_{i}}$, for $i=r+1,\ldots,r+l$, the linear functional $u_{i}^{*}$ depends only on the variables $z_{k_{r}+2i-r-1}$, $z_{k_{r}+2i-r}$ and for each $i\neq j$, $u_{j}^{*}$ does not depends on the variables $z_{k_{r}+2i-r-1}$ and $z_{k_{r}+2i-r}$. For $i=r+l+1$ to $k$, $u_{i}^{*}(z)=z_{i+t}$.
Define a projection $P:\mathbb{C}^{N}\to\mathbb{C}^{N}$ by
$P(z)=\sum_{j=1}^{k}u_{j}^{*}(z)u_{j}.$
Note that $P$ is a projection of norm $1$ onto $L$. Next we are looking for standard projections which commutes with $P$. For $i=1,\ldots,r$, consider the standard projection,
$\pi_{k_{i-1}+1,\ldots,k_{i}}(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$
Note that
$P\circ\pi_{k_{i-1}+1,\ldots,k_{i}}(z_{1},\ldots,z_{N})=P(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$
For each $i=1,\ldots,r$, note that $k_{i-1}+1,\ldots,k_{i}$ components of $P$ only depends on $z_{k_{i-1}+1},\ldots,z_{k_{i}}$. So, $\pi_{k_{i-1}+1,\ldots,k_{i}}\circ P(z_{1},\ldots,z_{N})=P(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$.
Let $m_{i}:=k_{r}+2i-r-1$. For $i=r+1,\ldots,r+l$, consider another projection
$\pi_{m_{i},m_{i}+1}(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{m_{i}-1},0,0,z_{m_{i+2}},\ldots,z_{N})$
Using similar arguments as mentioned above, we can conclude that $P\circ\pi_{m_{i},m_{i}+1}=\pi_{s_{i},s_{i}+1}\circ P$. As we noted before, the total number of dependent variables in the system of equation $AZ=0$ is $N-k+r+l$. So, we consider another projection $\pi$ defined by $\pi(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{s_{1}},0,\ldots,0)$. Note that $P$ commutes with $\pi$ and $rank(\pi)=N-k+r+l$. If $r+l\geq 1$, then rank of $\pi$ is atleast $N-(k-1)$. If $r+l=0$, then $rank(\pi)$ is atleast $(N-k)$. This finishes the proof of (i).
To prove (ii), take $P_{k}:=P$, which is a projection of norm one defined in the proof of (i). From the proof of (i), we obtain a standard projection $\pi_{k-1}$ of rank $(N-1)$ which commute with $P$. Let $P_{k-1}:=\pi_{k-1}\circ P$. Observe that $P_{k-1}$ is a norm one projection and $\rm image(P_{k-1})$ is a $k$-dimensional is a $(k-1)$-dimensional subspace of $\mathbb{C}^{N}$. Again applying the same argument to $P_{k-1}$, we get another projection $\pi_{k-2}$ of rank $(N-1)$ which commutes with $P$. Define $P_{k-2}:=\pi_{k-2}\circ P_{k-1}=\pi_{k-2}\circ\pi_{k-1}\circ P$. Hence we get that $P_{k-2}$ is a projection of norm one and it’s image is a $(k-2)$ dimensional subspace. Continuing this process, until we will obtain a projection $P_{1}$ of norm one such that $\rm image(P_{1})$ is a one dimensional subspace of $\mathbb{C}^{N}$. ∎
An apparently even more geometric formulation of the retracts of $\ell_{p}$ balls for $1\leq p<\infty,p\neq 2$ is that the retracts are precisely isometric copies of lower dimensional $\ell_{p}$-balls; see theorem 6.3 of [Beata-survey]. This is a feature shared by the $\ell_{p}$-‘balls’ even for positive values of the parameter $p<1$. Indeed, we now move onto this case dealt with by theorem 1.9. Before going into its proof 1.9, we need to make a couple of observations about the geometry of the $\ell_{q}$ ball $D_{q}$ in $\mathbb{C}^{N}$ (for $q<1$). Firstly, although its absolute image of is concave (easily and vividly seen in the two-dimensional case), $D_{q}$ is actually not concave (nor convex) domain in $\mathbb{C}^{N}$; not even locally, near any of the smooth boundary points. We may establish this by computing the real Hessian at any smooth boundary point with positive real coordinates as $D_{q}$ is multicircular; we show this in the two-dimensional setting for simplicity, while it holds in higher dimensions as well by straight-forward extension of the arguments and computations. Rewrite the defining function for $D_{q}$ as
$\varphi(z,w)=z^{q/2}\overline{z}^{q/2}+w^{q/2}\overline{w}^{q/2}-1$
To compute the real Hessian, note the following.
$H$ $=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial z\partial\overline{w}}\\ \frac{\partial^{2}\varphi}{\partial w\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial w\partial\overline{w}}\end{bmatrix}=\begin{bmatrix}\frac{q^{2}}{4}|z|^{q-2}&0\\ 0&\frac{q^{2}}{4}|w|^{q-2},\end{bmatrix}$
$\tilde{H}$ $=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z^{2}}&\frac{\partial^{2}\varphi}{\partial z\partial w}\\ \frac{\partial^{2}\varphi}{\partial w\partial z}&\frac{\partial^{2}\varphi}{\partial w^{2}}\end{bmatrix}=\begin{bmatrix}\frac{q}{2}(\frac{q}{2}-1)|z|^{q}z^{-2}&0\\ 0&\frac{q}{2}(\frac{q}{2}-1)|w|^{q}w^{-2}\end{bmatrix}$
Before proceeding, we remark that we shall utilize the above to compute the real Hessian only at the (smooth) absolute boundary points of $D_{q}$, such points being precisely those points $(x,y)$ of $\partial D_{q}$ both of whose coordinates are positive; while this suffices as $D_{q}$ is Reinhardt, it is important to note that the (tangent) vectors along which the Hessian form is evaluated need to be allowed to have non-real (complex) coordinates. Let $v=[v_{1}~{}v_{2}]$ and $\overline{v}=[\overline{v_{1}}~{}\overline{v_{2}}]$. Then the real Hessian at the point $(x,y)$ evaluated along the direction $v$ is
(5.1) $vH\overline{v}^{T}|_{(x,y)}+Re(vHv^{T})_{(x,y)}\\ =\frac{q^{2}}{4}\left(x^{q-2}|v_{1}|^{2}+y^{q-2}|v_{2}|^{2}\right)+\frac{q}{2}(\frac{q}{2}-1)Re\left(x^{q-2}v_{1}^{2}+y^{q-2}v_{2}^{2}\right).$
If $v_{1},v_{2}$ are purely imaginary, then the above expression (5.1) is strictly positive. If $v_{1}$ and $v_{2}$ are real, then the above expression (5.1) becomes
$\frac{q^{2}-q}{2}\left(x^{q-2}v_{1}^{2}+y^{q-2}v_{2}^{2}\right)$
which is negative as $q<1$. We conclude therefore that $D_{q}$ is neither convex nor concave (though of-course pseudoconvex) at $(x,y)\in\partial D_{q}$. Owing to already made remarks this holds at each point of $\partial D_{q}$ other than the negligible set of non-smooth boundary points of $D_{q}$.
@@
Proceeding now towards the retracts of $D_{q}$ (where we no longer assume $N=2$ for that would be an oversimplification unlike the matter in the preceding paragraph), it must first be noted that by proposition 1.4, every holomorphic retract of $\ell_{q}$-ball through the origin is actually linear even when $q<1$; for $1$-dimensional retracts, this follows from theorem 1.3 but as highlighted in the introduction, the thrust here is on retracts of arbitrary dimension. So, if $Y_{q}$ is a holomorphic retract of $D_{q}$ through the origin, then there exists a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$. Of-course, we are only interested in non-trivial retracts, so $Y_{q}$ is of the form $Y_{q}=D_{q}\cap Y$ for some $r$-dimensional $\mathbb{C}$-linear subspace $Y$ of $\mathbb{C}^{N}$. We only need to show that $Y$ cannot be of the form
(5.2) $\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:z_{i}=a_{1,i}z_{1}+\ldots+a_{r,i}z_{r}~{}\text{for}~{}r+1\leq i\leq N\},$
where not all $(a_{i,j})$ are zero.
@@
To prove this theorem 1.9 by contradiction as already indicated above, assume that $Y_{q}$ is a linear retract of $D_{q}$ and a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$. Let $\{e_{i}:1\leq i\leq N\}$ be the standard basis of $\mathbb{C}^{N}$. With notations as in (5.2), define another basis $\{v_{i}:1\leq i\leq N\}$ of $\mathbb{C}^{N}$, as follows:
@@
$t_{j}=\frac{1}{\left(1+\sum_{k=r+1}^{N}|a_{j,k}|^{q}\right)^{1/q}}.$
@@
$e_{j}=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}v_{l}\right).$
Applying $P$ on both sides, we get
$P(e_{j})=\frac{1}{t_{j}}P(v_{j})-\left(\sum_{l=r+1}^{N}a_{j,l}P(v_{l})\right).$
@@
(5.3) $P(v_{j})=\left(u_{1,j},\ldots,u_{r,j},\sum_{k=1}^{r}a_{k,r+1}u_{k,j},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,j}\right)~{}\text{for}~{}r+1\leq j\leq N.$
@@
$P(e_{j})=\frac{1}{t_{j}}v_{j}-\left(\sum_{l=r+1}^{N}a_{j,l}\left(u_{1,l},\ldots,u_{r,l},\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,\sum_{k=1}^{r}a_{k,N}u_{k,l}\right)\right)\\ =\left(-\sum_{l=r+1}^{N}a_{j,l}u_{1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{j-1,l},1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l},-\sum_{l=r+1}^{N}a_{j,l}u_{j+1,l},\ldots,-\sum_{l=r+1}^{N}a_{j,l}u_{r,l},\right.\\ \left.a_{j,r+1}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,r+1}u_{k,l},\ldots,a_{j,N}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,N}u_{k,l}\right).$
@@
(5.4) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}\left|a_{j,s}-\sum_{l=r+1}^{N}a_{j,l}\sum_{k=1}^{r}a_{k,s}u_{k,l}\right|^{q}\leq 1.$
Therefore, following inequality holds for $r+1\leq j\leq N$:
(5.5) $\sum_{k=1}^{r}|u_{k,j}|^{q}+\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}~{}u_{k,j}\right|^{q}\leq 1.$
By reverse triangle inequality for $q^{\text{th}}$ power where $10<q<$ 1: $|z-w|^{q}\geq\left||z|^{q}-|w|^{q}\right|$ for all $z,w\in\mathbb{C}$, inequality (5.4) becomes
$-\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0.$
@@
$-\sum_{l=r+1}^{N}|a_{j,l}|^{q}\left(1-\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}-\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}u_{k,j}\right|^{q}\right)+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}+\\ \sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0.$
Simplifying this inequality, we get
$\sum_{l=r+1}^{N}|a_{j,l}|^{q}\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}\leq 0.$
As all quantities on the left are obviously non-negative, it follows that
@@
Combining this result with inequality (5.4), we get
(5.7) $\left|1-\sum_{l=r+1}^{N}a_{j,l}u_{j,l}\right|\leq t_{j},\quad\text{ for each }j\in\{1,2,\ldots,r\}.$
By our hypothesis about the $a_{i,j}$’s, atleast one of $a_{i,j}$ is non-zero. If $a_{j_{0},l_{0}}\neq 0$ for some $j_{0},l_{0}$ then from (5.6) we obtain that $u_{j,l_{0}}=0$ for all $j$ with $j\neq j_{0}$. From (5.5), we get $|u_{j_{0},l_{0}}|\leq t_{j_{0}}$. Therefore, we conclude that for all $(j,l)$, we have $|u_{j,l}|\leq t_{j}$. Using (5.7), we get